OpenStudy (anonymous):

A table factor of .7312 from a present value table of $1 means that a certain rate of interest for a certain period of time it will equal?

OpenStudy (anonymous):

1/(1+r) = 0.7312 => 1+r = 1/0.7312 => r = 1/0.7312 - 1

OpenStudy (anonymous):

r = 0.3676 = 36.76%

OpenStudy (anonymous):

? * ? Did that help?

OpenStudy (anonymous):

kinda. i am hopeless at math.